On a sunny day, a building and its shadow form the sides of a right triangle. If the hypotenuse is 24 m long and the shadow is 15 m, how tall is the building?
d^2=24^2 + 15^2 = 576 +225 = 801
= 16.125
Could you tell me if this is correct?
Cynthia

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(57384)

On a sunny day, a building and its shadow form the sides of a right triangle. If the hypotenuse is 24 m long and the shadow is 15 m, how tall is the building?
d^2 = leg^2 + (other leg)^2
24^2 = x^2 + 15^2
---
x^2 = 24^2-15^2
x^2 = 351
x = sqrt(351) = 18.73 m
--------------------------------
Cheers,
Stan H.

Answer by richwmiller(9143)

Your math is all wrong. the square root of 801 has to be more that 20 since 20^2=400. sqrt(801)=28.3019
Also d has to be bigger than the other two.
Which brings us to the biggest problem.
The hypotenuse is the longest side of a right triangle.
You have the hypotenuse in the wrong place.
a^2+b^2=c^2
c is the hypotenuse (the slanted part) and a and b are the legs (up and across)
you have a and c
a=15 and c=24
a^2+b^2=c^2